OpenStudy (anonymous):

If g(x) + xsin(g(x)) = x^5 and g(4) = 0 , find g '(4)

OpenStudy (anonymous):

\[g(x)+x\sin(g(x))=x^5\] Differentiate both sides and solve for \(g'(x)\), you have \[\begin{align*}g'(x)+\sin(g(x))-x\cos(g(x))~g'(x)&=5x^4\\\\ g'(x)&=\frac{5x^4-\sin(g(x))}{1-x\cos(g(x))} \end{align*}\] Plug in \(x=4\) and make use of the given data points.